# Natural super-orbitals representation of many-body operators

### v1

@@ IV.1 Non-interacting systems
$\begin{split}\hat{\hat{q}}^{\dagger}_{n}&=(\sqrt{A}\hat{c}^{\dagger}_{n}+\hat{\tilde{c}}_{n}),\\ \hat{\hat{q}}^{\dagger}_{L+n}&=(\hat{c}^{\dagger}_{n}-\sqrt{A}\hat{\tilde{c}}_{n}),\\ \end{split}$ (30)
where $A=1-e^{-i\varepsilon_{n}t}$, such that $q^{\dagger}_{n}q^{\dagger}_{L+n}|0\rangle\!\rangle$ correctly reproduces Eq. (29). Note that these orbitals must be correctly normed. Each factorized operator $\hat{O}_{n}$ in Eq. (28) can thus be decomposed in two such kind of orbitals, which leads to:
@@ IV.2 Haar random operators
The case of random unitary operators is much simpler. It is possible to calculate the matrix elements $A_{mn}$ on average using the common machinery of Weingarten calculus [42], as shown in Appendix A. We find that for the time-evolution operator, these elements are all always zero, such that the spectrum of $R$ is fully degenerate with $\lambda=1/2$, corresponding to the worst case in which the maximal correlation entropy $S_{\mathrm{max}}=L\log(2)$ is reached. In such a case, there exist no one-particle basis in which a random unitary can be simplified. As for $\hat{U}(t)$ the eigenvalues of $A$ are simply zero, we also investigate the behavior of the natural super-orbitals for a local operator which is rotated by a random unitary. In this case, we show that the eigenvalues of the matrix $A$ are decaying with the system size as $2^{-L}$ which converges exponentially to a maximally correlated operator.
@@ V Results
We present the numerical study of the correlation spectrum for two specific unitary operators, the time-evolution operator $\hat{U}(t)$ and an evolved local operator of the type $(\hat{n}_{i}(t)-1/2)$ for two different models. The first model is the $t-V$ model, in which every site is interacting with their neighbors, and the second one is the interacting resonant level model (IRLM) in which the interaction is localized at an edge of a tight-binding chain. Our numerical simulations are based on a $U(1)$-symmetric MPO representation of the corresponding operators, and employ the time-evolving block decimation (TEBD) algorithm [43, 9] in real space with a maximum bond dimension $\chi=1024$, as well as a $4^{\mathrm{th}}$-order Trotter decomposition with a time step $dt=10^{-2}$ to ensure the convergence of the data up to the times shown in the plots.
@@ V.1 $t\text{-}V$ model
The first system investigated here is the $t-V$ model of fermions in a one-dimensional open chain. This model is equivalent to the Heisenberg spin chain (or $XXZ$ model) through a Jordan-Wigner transformation [44], and is studied here as a prototypical interacting system, in which interaction and hopping compete at every site. The Hamiltonian is the following:
@@
In the left panel of Fig. 1, we show the spectra of $R[\hat{U}(t)]$ and $R[\hat{n}_{i}(t)-1/2]$, where $i$ is a site chosen at the center of a chain of length $L=50$, for time evolution with parameters $\gamma=V=1$. The blue curves correspond to the unitary operator $\hat{U}(t)$ at different times, initially represented as an identity MPO, to which Trotter gates are sequentially applied on one side to simulate time evolution. At $t=0$, the spectrum reflects a trivial operator, consisting solely of eigenvalues 00 and $1$. As time progresses, the spectrum evolves towards that of a highly correlated operator, with most orbital occupations approaching $1/2$. The spectral gap $\Delta$—defined as the difference between the largest occupation among those close to 00 and the smallest among those close to $1$—initially equal to 1, closes over time in a non-algebraic manner. All orbitals become correlated similarly, lacking any notable internal structure except for the most and least occupied orbitals are affected by the open boundaries. The right panel of Fig. 1 presents the corresponding correlation entropy $S[\hat{U}(t)]$ as defined in Eq. (20), normalized by its maximal value $L\log(2)$. This entropy rapidly saturates, indicating that no preferential single-particle basis exists for $\hat{U}(t)$, and thus its structural complexity cannot be reduced via orbital rotations.
@@
The red curves illustrate the spectrum of $R[\hat{n}_{L/2}(t)-1/2]$ under the same conditions as for $\hat{U}(t)$. The initial operator is the identity at all sites except the center of the chain, where it is a local Pauli $S^{z}$ operator—thus also a trivial operator in terms of its natural super-orbitals. As time evolves, the spectrum deviates from the initial values 00 and $1$ in a structured manner: an increasing number of orbitals acquire occupations near $1/2$, while the rest remain close to 00 or $1$, with a smooth decay between the two regimes. This behavior reflects the light-cone structure characteristic of local operator spreading: at time $t$, the operator can be approximated as $\mathds{1}\otimes\cdots\otimes C\otimes\cdots\otimes\mathds{1}$, where $C$ denotes the non-trivial part of the operator within the light cone. If $C$ spans $c$ sites, then approximately $c$ orbitals exhibit occupations near $1/2$, while the remaining $L-c$ remain close to 00 or $1$. The smooth decay between these regions arises because the operator is not an exact tensor product between $C$ and the outside region. This structure also governs the behavior of the correlation entropy, which grows approximately linearly in time at a significantly slower rate than that of the time evolution operator. At late times, when the support of $C$ extends across the entire system, the entropy eventually saturates near its maximal value.
@@
A more detailed understanding of the structure of the correlation spectrum of $n_{i}(t)-1/2$ can be obtained by examining the spatial dispersion of the natural super-orbitals. In Fig. 2, we display the absolute values of the coefficients $v_{i\alpha}$, defined in Eq. (15), which describe the spatial profile of the natural orbital $\hat{q}_{\alpha}$ at time $t=8$. The orbital indices $\alpha$ are ordered by increasing proximity of their occupation numbers to 00 or $1$, such that the most correlated orbitals—those with occupations closest to 0.5—appear first. At this time, the light cone emerging from site $L=25$ has propagated approximately $8$ sites in both directions, which is reflected in the complex, delocalized structure of the leading natural super-orbitals for $\alpha\lesssim 16$. As the occupation numbers approach 00 or $1$, the orbitals become increasingly localized, consistent with the exponential decay of correlations between regions inside and outside the light cone [45, 17]. Ultimately, each natural orbital becomes localized on a single site, in agreement with the cartoon of the operator structure discussed above, wherein sites outside the light cone are effectively in a tensor product state. Note that the maximum amplitude of each orbital is $1/\sqrt{2}$, as the diagonalization of the matrix $R$ yields natural super-orbitals defined in the doubled Hilbert space, as introduced in Eq. (14). Only the first $L$ orbitals (out of a total of $2L$) are shown in Fig. 2, since the remaining orbitals are redundant.
@@ V.2 Impurity problem (IRLM)
Figure 3: Left: Correlation spectrum of the time evolution operator $\hat{U}(t)$ at different times for the IRLM, showing the half of orbitals the closest to 00, the other half being the particle-hole symmetric $1-\lambda_{\alpha}$. The plateau arising after $t\geq 8$ is due to the truncation error, as the corresponding MPO has reached the maximum bond dimension $\chi=1024$. The $x$-axis is cut at $\alpha=16$ instead of $\alpha=2L$ for visibility, as the remaining eigenvalues are below the machine precision. Right: Correlation entropy of $\hat{U}(t)$ as a function of time for different values of the interaction in the regime $U<D$, normalized by the maximal extensive value $S_{\mathrm{max}}=L\log{2}$. For both panels, we consider as system of $L=50$ sites. For both panels, $V=0.2D$.
@@
where the $1/2$ subtracted to the density operators enforce particle-hole symmetry, and $\gamma/2$ is taken for the half-bandwidth of the bath to be $D=1$ for $\gamma=1$. Hence we work in units of $D$ (or equivalently of $\gamma$) in the following. As for the $t-V$ model presented above, the correlation spectrum is symmetric under super-orbital rotations, which in this model also corresponds to flipping the signs of the parameters: $U\rightarrow-U$, as well as $V\rightarrow-V$ or $\gamma\rightarrow-\gamma$ yields the same properties of the correlation spectrum. In the IRLM, the ground state properties of the model with $U>0$ and $U<0$ are very different, as the Kondo temperature is exponentially smaller for negative values of the interaction, leading to a more correlated state [30]. Such an effect is not present here as the properties of the operator are related to its full spectrum, such that specific properties related to exceptional states (like the ground state) are completely diluted. In the following, we first discuss the properties of the time-evolution operator $\hat{U}(t)$ for different interaction strengths $0<U<D$ and $U>D$, and then we study the properties of an evolved local operator on the impurity site $\hat{n}_{1}(t)-1/2\equiv\hat{S}^{z}_{1}(t)$. The numerical parameters that we choose for our simulations are the same as the ones used for the Heisenberg model.
Time-evolution operator: The correlation spectrum of the operator $\hat{U}(t)$ is shown in the left panel of Fig. 3 for $U=0.3D$ in a system with $L=50$ sites. We present the spectrum at various times up to $t=9$, beyond which a plateau appears in the spectrum around $10^{-6}$, attributable to truncation errors. The figure displays only the part of the spectrum near 00, as the full spectrum is symmetric about $1$. The key observation is that, similar to the ground state of this model [30, 31], the occupations of the natural super-orbitals of the time-evolution operator decay exponentially at all times. However, this decay becomes progressively slower as time increases, eventually leading the spectrum to converge toward the maximal entropy value $S_{\mathrm{max}}=L\log(2)$.
@@
The right panel of Fig. 3 shows the time evolution of the correlation entropy for different interaction strengths in the weak-coupling regime $U<D$. After a timescale comparable to the electronic bandwidth $D$, the correlation entropy begins to grow linearly with time. As the interaction strength increases, the entropy grows more rapidly, indicating that the operator becomes more entangled—i.e., the occupations of a larger number of natural super-orbitals deviate more significantly from 00 and $1$. Nonetheless, the absolute value of the correlation entropy remains well below the maximal limit, implying that the operator admits a relatively compact representation in the natural orbital basis. In contrast, its real space MPO representation used in these simulations reaches the imposed bond dimension $\chi=1024$ already at $t=8$, due to the much faster growth of the von Neumann entropy in that basis. Rotation to natural super-orbitals thus offers great compression of the considered evolution operators.
@@
Eventually, the gap between the two groups vanishes, signaling a saturation in the complexity of the operator. This saturation is captured in the right panel of Fig. 5, which shows the evolution of the correlation entropy. The point of saturation appears to depend on the initial gap between the two orbital groups—related to the ratio $U/V$—and the rate of exponential decay within the first group, which is controlled by $U/D$.
@@ V.3 Early growth of the correlation entropy
Finally, we study the short time behaviour of the operators studied above for the two models presented here, the $t-V$ (or Heisenberg) model and the IRLM. At short times, the elements of the matrix $R$ of these unitary operators defined in Eq. (11) can be approximated using the following form of the Baker-Hausdorff-Campell (BHC) formula:
@@
Figure 8: Correlation entropy of the time-evolution operator $\hat{U}(t)$ (upper panel) and of an evolving local operator $\hat{n}_{i}-1/2$ (lower panel), for chains of $L=50$ in the $t-V$ model (blue curves) and in the IRLM model (red curves). The dashed curves indicate the expected scalings from perturbation theory (see text).
@@
For the Heisenberg model, the diagonal elements of $A$ are constant at leading order, with $A_{mm}=1/2$. For the off-diagonal elements, we find that $\hat{U}^{\dagger}\hat{c}_{m}^{\dagger}\hat{U}$ commutes with $\hat{c}_{n}$ up to order $p=|n-m|$ in the BCH expansion, implying $A_{mn}\propto t^{|n-m|}$. Consequently, retaining only the leading order contribution for $|n-m|=1$ reduces $A$ to a tridiagonal form. Following Eq. (13), we compute the spectrum of $A^{\dagger}A$ which is also a tridiagonal matrix. Its spectrum can then obtained analytically and is given by $\mu_{\alpha}=1/2-t^{2}c_{\alpha}/2$, where the coefficients $c_{\alpha}$ encode spectral details independent of $t$. Accordingly, the spectrum of $R[\hat{U}(t)]$ is given by $\lambda_{\alpha,+}=1-t^{2}c_{\alpha}$ and $\lambda_{\alpha,-}=t^{2}c_{\alpha}$. Due to the isotropy of the Heisenberg model, all eigenvalues exhibit the same scaling behavior, leading to an extensive number of natural super-orbitals becoming more correlated at the same time.
@@
In contrast, for the IRLM, the interaction anisotropy causes the diagonal elements of $A$ in the bath to vanish as $t^{|m|}$ for $m>2$, so that only the impurity contributions remain dominant at leading order. In this case, the scaling of eigenvalues remains $\lambda_{\alpha,+}=1-t^{2}c_{\alpha}$ and $\lambda_{\alpha,-}=t^{2}c_{\alpha}$, but applies only to the leading modes ($\alpha=1,2$), while all other eigenvalues are either 00 or $1$. These behaviors are illustrated in the upper panel of Fig. 8, which shows the correlation entropy for both models across different parameters for a system size of $L=50$. As predicted, the correlation entropy grows quadratically in time ($\propto t^{2}$) for both models. However, the amplitude is larger in the Heisenberg case, reflecting the fact that more natural super-orbitals contribute at the same leading order.
@@
The case of the local operator $\hat{n}_{i}(t)-1/2$ requires a slightly different treatment. As before, we set $\hat{X}=i\hat{H}t$, but now choose $\hat{Y}=\hat{n}_{i}$. For the $t-V$ (or Heisenberg) model, the site index $i$ is taken at the center of the chain ($i=L/2$), while for the IRLM, $i=0$ corresponds to the impurity site. In both models, the diagonal and off-diagonal elements of the correlation matrix decay faster than $1/t$ as their distance from site $i$ increases. Consequently, the dominant contributions at leading order arise from a reduced $(3\times 3)$ sub-matrix of $A$, corresponding to indices $m,n=i-1$ to $i+1$. The structural differences between the two Hamiltonians lead to distinct scaling behaviors of the spectrum of $A$: in the Heisenberg model, the spectrum scales as $S\propto t^{4}$ (first non-vanishing terms), whereas in the IRLM it scales as $S\propto t^{6}$. These predictions are confirmed by numerical simulations, as shown in the bottom panel of Fig. 8.
@@
VI Conclusion
In this work, we introduced the concept of natural super-orbitals associated with many-body operators, defined as the eigenvectors of the one-body super-density matrix constructed from a vectorized operator. To illustrate the meaning of these orbitals in the super-Hilbert space of $2L$ particles, we analyzed several simple cases. We demonstrated that a suitable super-orbital rotation can render an operator trivial in the MPO representation for non-interacting systems, while no such simplification is possible for random operators. Additionally, we showed that for unitary operators, a super-orbital rotation corresponds to a standard basis change that preserves the meaning of the operator in the physical Hilbert space of $L$ particles.
We then applied this framework to characterize the natural super-orbitals of both the time-evolution operator and a time-evolved local operator in two interacting models: the $t-V$ model and a quantum impurity model. In the impurity model, the natural super-orbitals reveal a compressed structure in both operators, evident through the exponential decay of their occupation numbers. This suggests that expressing these operators in the natural orbital basis allows for a significantly more compact MPO representation at all times, enabling the study of larger systems and longer time scales beyond the reach of current real-space tensor network methods. Using the correlation entropy, we observed that the complexity of a local operator initially located on the impurity site saturates over time, with the saturation timescale increasing with interaction strength. This behavior suggests that quantum correlations becomes absent at large scales, even at late times.

### v2

@@ IV.1 Non-interacting systems
[…]
where $A=1-e^{-i\varepsilon_{n}t}$, such that $\hat{\hat{q}}^{\dagger}_{n}\hat{\hat{q}}^{\dagger}_{L+n}|0\rangle\!\rangle$ correctly reproduces Eq. (30). Note that these orbitals must be correctly normed. Each factorized operator $\hat{O}_{n}$ in Eq. (29) can thus be decomposed in two such kind of orbitals, which leads to:
@@ IV.2 Haar random operators
The case of random unitary operators is much simpler. It is possible to calculate the matrix elements $A_{mn}$ on average using the common machinery of Weingarten calculus [66], as shown in Appendix A. We find that for the time-evolution operator, these elements are all always zero, such that the spectrum of $R$ is fully degenerate with $\lambda=1/2$, corresponding to the worst case in which the maximal correlation entropy $S_{\mathrm{max}}=L\log(2)$ is reached, indicating maximal non-Gaussianity and non-stabilizerness. In such a case, there exist no one-particle basis in which a random unitary can be simplified in terms of any of the resources that we presented. As for $\hat{U}(t)$ the eigenvalues of $A$ are simply zero, we also investigate the behavior of the natural super-orbitals for a local operator which is rotated by a random unitary. In this case, we show that the eigenvalues of the matrix $A$ are decaying with the system size as $2^{-L}$ which converges exponentially to a maximally correlated operator.
@@ V Results
We present the numerical study of the correlation spectrum for two specific unitary operators, the time-evolution operator $\hat{U}(t)$ and an evolved local operator of the type $(\hat{n}_{i}(t)-1/2)$ for two different models. The first model is the tVt-V model, in which every site is interacting with their neighbors, and the second one is the interacting resonant level model (IRLM) in which the interaction is localized at an edge of a tight-binding chain. Our numerical simulations are based on a $U(1)$-symmetric MPO representation of the corresponding operators, and employ the time-evolving block decimation (TEBD) algorithm [67, 9] in real space with a maximum bond dimension $\chi=1024$, as well as a $4^{\mathrm{th}}$-order Trotter decomposition with a time step $dt=10^{-2}$ to ensure the convergence of the data up to the times shown in the plots.
@@ V.1 $t\text{-}V$ model
The first system investigated here is the tVt-V model of fermions in a one-dimensional open chain. This model is equivalent to the Heisenberg spin chain (or $XXZ$ model) through a Jordan-Wigner transformation [68], and is studied here as a prototypical interacting system, in which interaction and hopping compete at every site. The Hamiltonian is the following:
@@
In the left panel of Fig. 1, we show the spectra of $R[\hat{U}(t)]$ and $R[\hat{n}_{i}(t)-1/2]$, where $i$ is a site chosen at the center of a chain of length $L=50$, for time evolution with parameters $\gamma=V=1$. The blue curves correspond to the unitary operator $\hat{U}(t)$ at different times, initially represented as an identity MPO, to which Trotter gates are sequentially applied on one side to simulate time evolution. At $t=0$, the spectrum reflects a trivial operator, consisting solely of eigenvalues 0 and $1$. As time progresses, the spectrum evolves towards that of a highly correlated operator, with most orbital occupations approaching $1/2$. The spectral gap $\Delta$—defined as the difference between the largest occupation among those close to 0 and the smallest among those close to 11—initially equal to 1, closes over time in a non-algebraic manner. All orbitals become correlated similarly, lacking any notable internal structure except for the most and least occupied orbitals are affected by the open boundaries. The right panel of Fig. 1 presents the corresponding correlation entropy $S[\hat{U}(t)]$ as defined in Eq. (25), normalized by its maximal value $L\log(2)$. This entropy rapidly saturates, indicating that no preferential single-particle basis exists for $\hat{U}(t)$, and thus its structural complexity cannot be reduced via orbital rotations.
@@
The red curves illustrate the spectrum of $R[\hat{n}_{L/2}(t)-1/2]$ under the same conditions as for $\hat{U}(t)$. The initial operator is the identity at all sites except the center of the chain, where it is a local Pauli $S^{z}$ operator—thus also a trivial operator in terms of its natural super-orbitals. As time evolves, the spectrum deviates from the initial values 0 and $1$ in a structured manner: an increasing number of orbitals acquire occupations near $1/2$, while the rest remain close to 0 or $1$, with a smooth decay between the two regimes. This behavior reflects the light-cone structure characteristic of local operator spreading: at time $t$, the operator can be approximated as $\mathds{1}\otimes\cdots\otimes C\otimes\cdots\otimes\mathds{1}$, where $C$ denotes the non-trivial part of the operator within the light cone. If $C$ spans $c$ sites, then approximately $c$ orbitals exhibit occupations near $1/2$, while the remaining LcL-c remain close to 0 or $1$. The smooth decay between these regions arises because the operator is not an exact tensor product between $C$ and the outside region. This structure also governs the behavior of the correlation entropy, which grows approximately linearly in time at a significantly slower rate than that of the time evolution operator. It also bounds the growth of non-stabilizerness for such a local operator, which has been show to scale linearly in time [58] due to the factorized form, as explained earlier. At late times, when the support of $C$ extends across the entire system, the entropy eventually saturates near its maximal value.
@@
A more detailed understanding of the structure of the correlation spectrum of $\hat{n}_{i}(t)-1/2$ can be obtained by examining the spatial dispersion of the natural super-orbitals. In Fig. 2, we display the absolute values of the coefficients $v_{i\alpha}$, defined in Eq. (15), which describe the spatial profile of the natural orbital $\hat{q}_{\alpha}$ at time $t=8$. The orbital indices $\alpha$ are ordered by increasing proximity of their occupation numbers to 0 or $1$, such that the most correlated orbitals—those with occupations closest to 0.5—appear first. At this time, the light cone emerging from site $L=25$ has propagated approximately $8$ sites in both directions, which is reflected in the complex, delocalized structure of the leading natural super-orbitals for $\alpha\lesssim 16$. As the occupation numbers approach 0 or $1$, the orbitals become increasingly localized, consistent with the exponential decay of correlations between regions inside and outside the light cone [69, 17]. Ultimately, each natural orbital becomes localized on a single site, in agreement with the cartoon of the operator structure discussed above, wherein sites outside the light cone are effectively in a tensor product state. Note that the maximum amplitude of each orbital is $1/\sqrt{2}$, as the diagonalization of the matrix $R$ yields natural super-orbitals defined in the doubled Hilbert space, as introduced in Eq. (14). Only the first $L$ orbitals (out of a total of $2L$) are shown in Fig. 2, since the remaining orbitals are redundant.
@@ V.2 Impurity problem (IRLM)
where the $1/2$ subtracted to the density operators enforce particle-hole symmetry, and $\gamma/2$ is taken for the half-bandwidth of the bath to be $D=1$ for $\gamma=1$. Hence we work in units of $D$ (or equivalently of $\gamma$) in the following. As for the tVt-V model presented above, the correlation spectrum is symmetric under super-orbital rotations, which in this model also corresponds to flipping the signs of the parameters: $U\rightarrow-U$, as well as $V\rightarrow-V$ or $\gamma\rightarrow-\gamma$ yields the same properties of the correlation spectrum. In the IRLM, the ground state properties of the model with $U>0$ and $U<0$ are very different, as the Kondo temperature is exponentially smaller for negative values of the interaction, leading to a more correlated state [30]. Such an effect is not present here as the properties of the operator are related to its full spectrum, such that specific properties related to exceptional states (like the ground state) are completely diluted. In the following, we first discuss the properties of the time-evolution operator $\hat{U}(t)$ for different interaction strengths $0<U<D$ and $U>D$, and then we study the properties of an evolved local operator on the impurity site $\hat{n}_{1}(t)-1/2\equiv\hat{S}^{z}_{1}(t)$. The numerical parameters that we choose for our simulations are the same as the ones used for the Heisenberg model.
Figure 3: Left: Correlation spectrum of the time evolution operator $\hat{U}(t)$ at different times for the IRLM, showing the half of orbitals the closest to 0, the other half being the particle-hole symmetric $1-\lambda_{\alpha}$. The plateau arising after $t\geq 8$ is due to the truncation error, as the corresponding MPO has reached the maximum bond dimension $\chi=1024$. The $x$-axis is cut at $\alpha=16$ instead of $\alpha=2L$ for visibility, as the remaining eigenvalues are below the machine precision. Right: Correlation entropy of $\hat{U}(t)$ as a function of time for different values of the interaction in the regime $U<D$, normalized by the maximal extensive value $S_{\mathrm{max}}=L\log{2}$. For both panels, we consider as system of $L=50$ sites. For both panels, $V=0.2D$.
Time-evolution operator: The correlation spectrum of the operator $\hat{U}(t)$ is shown in the left panel of Fig. 3 for $U=0.3D$ in a system with $L=50$ sites. We present the spectrum at various times up to $t=9$, beyond which a plateau appears in the spectrum around $10^{-6}$, attributable to truncation errors. The figure displays only the part of the spectrum near 0, as the full spectrum is symmetric about $1$. The key observation is that, similar to the ground state of this model [30, 31], the occupations of the natural super-orbitals of the time-evolution operator decay exponentially at all times. However, this decay becomes progressively slower as time increases, eventually leading the spectrum to converge toward the maximal value of the correlation entropy $S_{\mathrm{max}}=L\log(2)$.
@@
The right panel of Fig. 3 shows the time evolution of the correlation entropy for different interaction strengths in the weak-coupling regime $U<D$. After a timescale comparable to the electronic bandwidth $D$, the correlation entropy begins to grow linearly with time. As the interaction strength increases, the entropy grows more rapidly, indicating that the operator becomes more entangled—i.e., the occupations of a larger number of natural super-orbitals deviate more significantly from 0 and $1$. Nonetheless, the absolute value of the correlation entropy remains well below the maximal limit, implying that the operator admits a relatively compact representation in the natural orbital basis. In contrast, its real space MPO representation used in these simulations reaches the imposed bond dimension $\chi=1024$ already at $t=8$, due to the much faster growth of the von Neumann entropy in that basis. Rotation to natural super-orbitals thus offers great compression of the considered evolution operators.
The linear growth of $S_{\mathrm{corr}}$ indicates that the non-Gaussianity of $\hat{U}(t)$ increases steadily over time. Additionally, the correlation spectrum allows us to estimate an upper bound on the non-stabilizerness generated by time evolution under a quantum impurity Hamiltonian. At any given time, the occupations of the natural super-orbitals decay exponentially. Therefore, orbitals with occupations below a given threshold $\delta$ can be neglected, as they effectively factorize and do not contribute to the measure of non-stabilizerness. We observe that the number of orbitals with occupations $n_{\alpha}>\delta$ increases linearly with time as $2vt$, where the factor $2$ comes from the almost occupied and empty orbitals coming together and $v$ is a parameter dependent factor, which is larger for larger interaction $U$. Consequently, the number of Pauli strings with non-negligible coefficients (for arbitrarily small $\delta$) in the natural super-orbital basis grows proportionally to $c2^{2vt}$, where $c$ is a constant depending on system parameters and $\delta$. In the worst-case scenario, where all Pauli strings contribute equally with $a_{S}=1$, the non-stabilizerness of the time-evolved operator is bounded by:
$\mathcal{M}_{\mathrm{NSO}}(\hat{U}(t))\leq 2vt\log(2)+\tilde{c}$, (35)
where $\tilde{c}$ is a constant and the subscript NSO refers to the natural super-orbital basis. This bound can be further tightened by accounting for the exponential decay of super-orbital occupations above the threshold $\delta$, which imposes additional structure on the coefficients $a_{S}$. Equation (35) resembles the known bound on the growth of magic for local operators, which also scales linearly in time [58] due to the factorization of real-space orbitals outside the light cone. It is somewhat surprising that the time-evolution operator $\hat{U}(t)$ displays a light-cone-like structure, given that it is completely non-local in real space. However, in the natural super-orbital basis, the majority of orbitals factorize due to the locality of the underlying interaction and thus do not contribute significantly to the Pauli decomposition. As time evolves, local interactions cause correlations to spread across more orbitals, increasing the number of non-negligible Pauli strings in a manner reminiscent of a light-cone structure, despite the operator’s inherent non-locality.
@@
Eventually, the gap between the two groups vanishes, signaling a saturation in the complexity of the operator. This saturation is captured in the right panel of Fig. 5, which shows the evolution of the correlation entropy. The point of saturation appears to depend on the initial gap between the two orbital groups—related to the ratio U/VU/V—and the rate of exponential decay within the first group, which is controlled by $U/D$.
@@ V.3 Early growth of the correlation entropy
Finally, we study the short time behaviour of the operators studied above for the two models presented here, the tVt-V (or Heisenberg) model and the IRLM. At short times, the elements of the matrix $R$ of these unitary operators defined in Eq. (11) can be approximated using the following form of the Baker-Hausdorff-Campell (BHC) formula:
@@
Figure 8: Correlation entropy of the time-evolution operator $\hat{U}(t)$ (upper panel) and of an evolving local operator $\hat{n}_{i}-1/2$ (lower panel), for chains of $L=50$ in the tVt-V model (blue curves) and in the IRLM model (red curves). The dashed curves indicate the expected scalings from perturbation theory (see text).
@@
For the Heisenberg model, the diagonal elements of $A$ are constant at leading order, with $A_{mm}=1/2$. For the off-diagonal elements, we find that $\hat{U}^{\dagger}\hat{c}_{m}^{\dagger}\hat{U}$ commutes with $\hat{c}_{n}$ up to order p=|nm|p=|n-m| in the BCH expansion, implying $A_{mn}\propto t^{|n-m|}$. Consequently, retaining only the leading order contribution for |nm|=1|n-m|=1 reduces $A$ to a tridiagonal form. Following Eq. (13), we compute the spectrum of $A^{\dagger}A$ which is also a tridiagonal matrix. Its spectrum can then obtained analytically and is given by $\mu_{\alpha}=1/2-t^{2}c_{\alpha}/2$, where the coefficients $c_{\alpha}$ encode spectral details independent of $t$. Accordingly, the spectrum of $R[\hat{U}(t)]$ is given by $\lambda_{\alpha,+}=1-t^{2}c_{\alpha}$ and $\lambda_{\alpha,-}=t^{2}c_{\alpha}$. Due to the isotropy of the Heisenberg model, all eigenvalues exhibit the same scaling behavior, leading to an extensive number of natural super-orbitals becoming more correlated at the same time.
@@
In contrast, for the IRLM, the interaction anisotropy causes the diagonal elements of $A$ in the bath to vanish as $t^{|m|}$ for $m>2$, so that only the impurity contributions remain dominant at leading order. In this case, the scaling of eigenvalues remains $\lambda_{\alpha,+}=1-t^{2}c_{\alpha}$ and $\lambda_{\alpha,-}=t^{2}c_{\alpha}$, but applies only to the leading modes ($\alpha=1,2$), while all other eigenvalues are either 0 or $1$. These behaviors are illustrated in the upper panel of Fig. 8, which shows the correlation entropy for both models across different parameters for a system size of $L=50$. As predicted, the correlation entropy grows quadratically in time ($\propto t^{2}$) for both models. However, the amplitude is larger in the Heisenberg case, reflecting the fact that more natural super-orbitals contribute at the same leading order.
@@
The case of the local operator $\hat{n}_{i}(t)-1/2$ requires a slightly different treatment. As before, we set $\hat{X}=i\hat{H}t$, but now choose $\hat{Y}=\hat{n}_{i}$. For the tVt-V (or Heisenberg) model, the site index $i$ is taken at the center of the chain ($i=L/2$), while for the IRLM, $i=1$ corresponds to the impurity site. In both models, the diagonal and off-diagonal elements of the correlation matrix decay faster than $1/t$ as their distance from site $i$ increases. Consequently, the dominant contributions at leading order arise from a reduced $(3\times 3)$ sub-matrix of $A$, corresponding to indices m,n=i1m,n=i-1 to $i+1$. The structural differences between the two Hamiltonians lead to distinct scaling behaviors of the spectrum of $A$: in the Heisenberg model, the spectrum scales as $S_{\mathrm{corr}}\propto t^{4}$ (first non-vanishing terms), whereas in the IRLM it scales as $S_{\mathrm{corr}}\propto t^{6}$. These predictions are confirmed by numerical simulations, as shown in the bottom panel of Fig. 8.
@@
VI Conclusion and outlook
In this work, we introduced the concept of natural super-orbitals associated with many-body operators, defined as the eigenvectors of the one-body super-density matrix constructed from a vectorized operator. To illustrate the meaning of these orbitals in the super-Hilbert space of $2L$ particles, we analyzed several simple cases. We demonstrated that a suitable super-orbital rotation can render an operator trivial in the MPO representation for non-interacting systems, while no such simplification is possible for random operators. We introduced measures of non-stabilizerness and non-Gaussianity for operators and discuss their properties related to natural super-orbitals. Additionally, we showed that for unitary operators, a super-orbital rotation corresponds to a standard basis change that preserves the meaning of the operator in the physical Hilbert space of $L$ particles.
We then applied this framework to characterize the natural super-orbitals of both the time-evolution operator and a time-evolved local operator in two interacting models: the tVt-V model and a quantum impurity model. In the impurity model, the natural super-orbitals reveal a compressed structure in both operators, evident through the exponential decay of their occupation numbers. This suggests that expressing these operators in the natural orbital basis allows for a significantly more compact MPO representation at all times, enabling the study of larger systems and longer time scales beyond the reach of current real-space tensor network methods. We also discussed the implication of the exponential decay in terms of quantum resources, and showed that the non-stabilizerness of the time evolution operator of the impurity model has light-cone-like properties in the natural super-orbital basis stemming from the locality of the interaction. Using the correlation entropy, we observed that the complexity or non-Gaussianity of a local operator initially located on the impurity site saturates over time, with the saturation timescale increasing with interaction strength. This behavior suggests that quantum correlations becomes absent at large scales, even at late times.